OpenStudy (anonymous):

find the radius of the sphere with the given volume volume=32/3 pi ft^3

OpenStudy (anonymous):

24 inches

OpenStudy (nincompoop):

what is the formula for volume of a sphere?

OpenStudy (anonymous):

I dont know?

OpenStudy (nincompoop):

why don't you look it up?

OpenStudy (anonymous):

I will

OpenStudy (anonymous):

\[\frac{4}{3}\pi r^3 = \frac{32}{3} \pi _{ft^3}\] Divide by sides by \[\pi\] \[\frac{4}{3}r^3 = \frac{32}{3}_{ft^3}\] Divide both sides by \[\frac{4}{3}\] \[r^3 = \frac{\frac{32}{3}}{\frac{4}{3}}\] Simplify: \[r^3 = \frac{32}{3}\times\frac{3}{4}\] \[r^3 = \frac{96}{12}\] \[r^3 = 8\] What number multiplied by itself three times is 8? That number is your answer.